OpenStudy (anonymous):

i'll fan+i'll give medal!! Select the pair of equations whose graphs are perpendicular. A. 2x – 8y = 9 12x – 3y = 7 B. 3x + 6y = 8 y = 2x – 8 C. y = 2x – 7 x + y = 3 D. 2y = –3x + 5 2x + 3y = 4

OpenStudy (smileyxl3):

2y = -3x + 5 Gradient = -3/2 2x + 3y = 4 Gradient = -2/3 -1/(-2/3) = -3/-2 = 3/2 not equal to -3/2 so not perpendicular.

OpenStudy (anonymous):

First find the standard form of the equations. When you do that, you should get b as your answer.

OpenStudy (anonymous):

@smileyxl3 , yours is a good answer, but I think gradient is not suitable for her :). I think the idea of slope is better. Slope of a line * Slope of the other line = -1 with y = mx + b, slope is m. so @xellaphantx you can find m and answer for yourself

OpenStudy (anonymous):

I meant b as in the answer choices.

OpenStudy (smileyxl3):

slope of:3x + 6y = 8 6y = -3x+8 y = (-1/2)x+8/6 (slope is -1/2). slope of: y = 2x - 8 (slope is 2) (-1/2)(2) = -1 this is perpendicular